# A proof of the Kim-Vu sandwich conjecture

### v1

@@ 1 Introduction
We outline our methods in context with previous work in Section 2, but for now make some brief remarks. As in [13, 14, 15], we work via switching techniques, requiring the counting of certain ‘switching paths’. The accuracy we use for our bounds has to be carefully chosen, particularly in the most critical regime, and our analysis is completely separate to previous work. Very broadly, our improvements come in two main areas. Firstly, we work much more directly with random graphs than [13], which used pseudorandom properties from the spectral properties of random graphs. Secondly, our bounds on the number of switching paths do not hold for every $d$-regular subgraph of our random graph (as they do in [13]), but typically hold for sufficiently many for our proof. This allows us to pivot in a slight but very useful way, considering instead ‘switching paths’ between two random graphs rather than between a random graph and any one of its $d$-regular subgraphs.
@@ 2.1 Notation
All logarithms are natural. We use ‘hierarchy’ notation to manage the constants we use, where $\alpha\ll\beta$ means that there is some non-negative decreasing function $f$ such that what follows will hold for all $\alpha\leq f(\beta)$. For hierarchies with more than 2 variables these functions are chosen from right to left. For any $a,b,c\in\mathbb{R}$, we say $a=b\pm c$ if $b-c\leq a\leq b+c$.
@@ 2.2.1 Coupling processes
While doing this, we will simultaneously build a random graph $G_{*}$, starting similarly with the graph with vertex set $[n]$ and no edges. When we consider the random edge $e$, we also consider whether to add the random edge $e$ to $G_{*}$ or not; when $e\in E(G_{*})$ we do this with the fixed probability $1-\eta$, where $0<\eta\ll\varepsilon$. Then, if $e$ is not in $F$ when it is considered and the probability at (1) is at least $1-\eta$, we can couple the random addition of $e$ so that $e$ is always added to $F$ if it is added to $G_{*}$. As discussed below, with high probability, for most of this process, the probability at (1) is very close to 1. Thus, if we build such a random graph $G_{*}$ and stop when it has $\lceil(1-2\eta)dn/2\rceil$ edges, then with high probability we will have $G_{*}\subset F$.
@@
This, however, is not good enough. Our ‘error term’ $\eta$ in (4) will need to change during the process. For some $C_{0},\mu$ with $1/n\ll 1/C_{0}\ll\mu\ll\varepsilon$, we will set $\eta_{m}=C_{0}\max\Big\{\frac{\mu}{\log n},\left(\frac{n\log n}{m}\right)^{1/8}\Big\}$, for each $m$, and our desired value of $\eta$ in (4) will be (roughly) $1-\eta_{e(F)}$, so that we can show that (3) is always at least the right hand side of (5) with (roughly) $\eta_{e(F)}$ replacing $\eta$. From (5) (and as proved in Section 3.5.5), as we will use this for each value of $e(F)$ from $\binom{n}{2}$ down to $\left\lfloor\left(1+\frac{\varepsilon}{2}\right)\frac{dn}{2}\right\rfloor$, one key point is that the sum of $\frac{\eta_{m}dn/2}{m-dn/2}$ over $m=e(F)$ in this range is $O(C_{0}\mu/\varepsilon)$, so that this will be much smaller than $\varepsilon$.
@@
Like Lemma 2.2, we will prove Theorem 2.4 using switching methods, but this will be very significantly more difficult. Why this is more challenging is explained below, but the new ideas we use to do this are the heart of our ability to analyse this process as long as $d=\omega(\log n)$, compared to the regime $d\gg n/\sqrt{\log n}$ in which this was done by Gao, Isaev and McKay [15]. (In [13] a modified two-step process is used to neatly avoid some of these difficulties.)
@@
To use the switching method, we define a notion of ‘switching’ and consider when we can transform $K\in\mathcal{K}^{e}$ into $K^{\prime}\in\mathcal{K}^{f}$ using a switching. Comparing the number of ways we can do such a transformation on $K$ with the number of ways we could transform into $K^{\prime}$, will allow us to compare $|\mathcal{K}^{e}|$ to $|\mathcal{K}^{f}|$. This is best seen via an example.
@@
and, as $\eta\ll 1$, we thus have $|\mathcal{K}^{e}|\leq(1-\eta)|\mathcal{K}^{f}|$.
Letting $K\in\mathcal{K}^{e}$, for each $K^{\prime}\in\mathcal{K}^{f}$ with $KK^{\prime}\in E(L)$, $E(K)\triangle E(K^{\prime})$ is the edge set of a length 10 cycle $S$ which alternates between $K\setminus F$ and $K_{n}\setminus K$ with $e$ and $f$ on opposite sides. Thus, we can bound $d_{L}(K)$ by counting such cycles. We will have that the vertex degrees of $F$ are all $(1\pm\eta/100)\Delta$ for some $\Delta$ (in Lemma 2.2 this will be d2m/nd-2m/n). Arbitrarily labelling $V(e)=\{v_{1},v_{6}\}$, we will use the notation in Figure 1 to choose such a cycle $S$ vertex by vertex. After labelling the vertices of $f$ as $v_{5}$ and $v_{10}$ (with 2 choices in total), and then choosing $v_{2}$ and $v_{7}$ (with at most $n$ choices for each), we then have at most $(1+\eta/100)\Delta$ choices for choosing each of $v_{3},v_{4},v_{8},v_{9}$, as they are the neighbour in $K\setminus F$ of some previously chosen vertex. Therefore, we have $d_{L}(K)\leq(1+\eta/100)^{4}\Delta^{4}n^{2}\leq(1+\eta/3)\Delta^{4}n^{2}$.
@@
For Lemma 2.2, as $F\sim F^{-}(n,d,m)$, with $m=\omega(n\log n)$ and $d=\omega(\log n)$, a simple concentration inequality (see Section 2.3) and a union bound shows that $F\setminus K$ will be approximately regular, allowing this sketch to be carried out when $d=o(n)$. For larger $d$ it is harder to count these switching cycles as accurately, as no longer is $K_{n}\setminus K$ so dense. In previous work [13], inequalities like (2) in such a regime were shown to be likely to hold using complex analytic methods. Here, we will use instead two simple ideas to complete the proof when $d=\Omega(n)$. These ideas are very useful for the upper part of the sandwich, so we will introduce them now in the simpler setting of the lower part.
@@
Then, in order to show it is likely that most of the graphs $K\in\mathcal{K}_{e}\cup\mathcal{K}_{f}$ satisfy some pseudorandom conditions, we exploit the distribution of $F\sim F^{-}(n,d,m)$. Intuitively, if we form $F$ by taking $K\sim K_{d}(n)$ and deleting $m$ random edges, then if a property is very likely to hold for $K$ it should be likely to hold for most of the $d$-regular subgraphs containing $F$. (This is proved with the exact parameters we use in Sections 9.1 and 10.5.) This means that instead of studying the properties of the $d$-regular subgraphs containing $F$, we can more simply study a random $d$-regular graph.
Switching methods for the upper part of the sandwich, i.e., for Theorem 2.4. After using the two ideas above, carrying out a similar switching argument for the upper part (i.e., for Theorem 2.4) will essentially require us to show that if $K\sim K_{d}(n)$ and $F$ is formed from $K$ by adding $m$ edges uniformly at random from $n]^{(2)}\setminus E(K)$ (and thus $F\sim F(n,d,m)$) then with high probability we have the following for some suitable $\ell$.
@@
Compared to similar tasks in the lower part, this is much more difficult in the region $d=o(n)$ as no longer is one of these graphs very close to being complete. Our previous discussion would correspond to paths of length $2\ell=4$, but now we will have to take $\ell=\Omega(\log n/\log\log n)$ when $d=\log^{O(1)}n$, otherwise between some distinct $x,y\in[n]$ no such alternating paths will exist. This pinpoints why the process for the upper part is so much harder to analyse via switchings than the process for the lower part. In their work, Gao, Isaev, and McKay [13] used two processes, one run after the other, so that (very roughly), when such $(F\setminus K,K)$-alternating paths were found, at least $K$ was sparse when compared to $F\setminus K$ (see [13] for more details).
When $F\setminus K$ is sufficiently dense, with $m\geq n^{1+\sigma}$ for any fixed $\sigma>0$, we use results of Kim and Vu [19] on the concentration of multivariate polynomials to get sufficiently good bounds on the counts for A (see Section 8.2). Thus, we can fix some small $\sigma>0$, and assume that $d\leq n^{\sigma}$ and $m\leq n^{1+\sigma}$. Take distinct $x,y\in[n]$ and let $\delta=2m/n$, and note that we may expect FKF-K to have vertex degrees $(1+o(1))\delta$. Due to the randomness of $F\setminus K$ and $K$, if we walk out randomly from $x$ using edges alternately from FKF-K and then $K$ we should expect this random walk to rapidly mix so that, as long as (say) $\ell\geq 2\log n/\log(\delta d)$, after $2\ell$ edges we are roughly no more likely to arrive at $y$ than any other vertex. Heuristically, considering the vertex degrees in $F\setminus K$ and in the $d$-regular graph $K$, we should expect there to be likely to be $(1+o(1))\delta^{\ell}d^{\ell}/n$ $(F\setminus K,K)$-alternating $x,y$-paths of length $2\ell$.
Bounds of this form were proved in certain regions (roughly when $\delta\gg d\log n$) by Gao, Isaev and McKay [13] by working via the spectral properties of binomial random graphs (and using only that $K$ is $d$-regular), while similar counts on paths within a single graph have been shown by showing certain Markov chains mix rapidly (see, for example, as in [18], building on work in [8]). For Theorem 2.4, we need to work with the two different random graph models, in particular using properties of the more difficult to study random regular graph, and we need to deal with various subtleties due to the accuracy we need and the weaker bound on $d$ that we use. As far as is convenient, then, we rely on pseudorandom conditions of the random graphs involved, in part so that the results we prove may ultimately perhaps prove useful for the study of Conjecture 1.3.
@@
Roughly, then, we count walks alternating between $F\setminus K$ and $K$ using these sums of degrees, and show that most of these walks are actually paths. To do so we rely on pseudorandom properties that later appear as LABEL:eq:fewdoubled:K and LABEL:eq:fewdoubled:F-K. Then, considering alternating paths from $x$ and $y$ respectively to sets of different endvertices we use lower bounds on the number of edges in FKF-K between these different endvertices (using LABEL:prop:F-Kconnection later)to give a lower bound between the number of alternating paths between $x$ and $y$. A subtlety that is hard to anticipate from our discussion so far is that perhaps the most difficult case for Theorem 2.4 is when $d=C\log n$ and $\delta=\log^{O(1)}n$, but $\delta=\omega(\log^{8}n)$. This is because the error term, essentially $\eta$, that we wish to use in the counting of our alternating paths is to be smaller than $C_{0}\mu/\log n$ (see (6)), but we still have $\ell=\Omega(\log n/\log\log n)$, so that when we build our alternating paths of length $\ell$ from $x$, if the next edge from the current endpoint, $v$ say, is to be in $K$ then $v$ may have a fraction of neighbours in the path so far which exceeds $\eta d$, which here is $O(1)$. This creates some additional small but important complication that, for example, is reflected in the error term we use in LABEL:prop:lowerboundKFKalternating:newnewnew in Theorem 6.1 later.
To show that the number of $(F\setminus K,K)$-alternating $x,y$-paths of length $2\ell$ is likely bounded above, we exploit our likely lower bound (applied with a slightly different path length) as well as an inductive upper bound on the number of alternating paths from $y$. Revealing the edges in $F\setminus K$ not containing $x$, we deduce that most vertices in $n]\setminus\{x,y\}$ – those not in $Z$, say – have at most $(1+O(\eta))\delta^{\ell-1}d^{\ell}/n$ $(K,F\setminus K)$-alternating paths of length $2\ell-1$ to $y$. Then, revealing the edges in $F\setminus K$ next to $x$, we get that there are at most $(1+O(\eta))\delta^{\ell}d^{\ell}/n$ $(K,F\setminus K)$-alternating $x,y$-paths of length $2\ell$ which do not have their second vertex in $Z\cap N_{K}(x)$. As $Z$ was a relatively small set (it will have $O(\eta n)$ vertices), using this and similar properties iteratively will allow us to give an upper bound for all the $x,y$-paths we are interested in apart from those in an increasingly constrained set of paths. After sufficiently many iterations, a simple bound on the remaining possibilities will be sufficient to give us our desired upper bound. This argument is carried out in Section 7.
@@ 2.2.3 Paper outline
In Section 3, we give the details of the edge deletion process we use to create the upper part of the sandwich, reducing the proof of the upper part of the sandwich to Theorem 2.4. In Section 4, using our main switching argument we reduce proving Theorem 4.1 to proving certain pseudorandom conditions concerning switching paths are likely to hold for random graphs $F$ and $K$. In Sections 57 we prove these pseudorandom conditions are likely to hold when $m\geq n^{1+\sigma}$ (for any fixed $\sigma>0$) In Section 5, we give likely upper bounds on the number of $(F\setminus K,K)$-alternating paths of a given length which start from a fixed vertex. In Section 6, we give likely lower bounds on the number of $(F\setminus K,K)$-alternating paths between any two fixed vertices (moreover, on such paths which avoid an arbitrary small set of vertices). In Section 7, we give likely upper bounds on the number of $(F\setminus K,K)$-alternating paths between any two fixed vertices. In Section 8, we give similar bounds to Sections 57 when $m\geq n^{1+\sigma}$ (for any fixed small $\sigma$). In Section 9, we put this together to prove Theorem 2.4. In combination, this completes the proof for the upper part of the sandwich. In Section 10, we give the lower part of the sandwich and its analysis.

### v2

@@ 1 Introduction
We outline our methods in context with previous work in Section 2, but for now make some brief remarks. As in [13, 14, 15], we work via switching techniques, requiring the counting of certain ‘switching paths’. The accuracy we use for our bounds has to be carefully chosen, particularly in the most critical regime, and our analysis is completely separate to previous work. Very broadly, our improvements come in two main areas. Firstly, we work much more directly with random graphs than [13], which used pseudorandom properties from the spectral properties of random graphs. Secondly, our bounds on the number of switching paths do not hold for every $d$-regular subgraph of our random graph (as they do in [13]), but typically hold for sufficiently many for our proof. This allows us to pivot in a slight but very useful way, considering instead ‘switching paths’ between two random graphs (a random $d$-regular graph and a random graph containing it) rather than between a random graph and any one of its $d$-regular subgraphs.
@@ 2.1 Notation
All logarithms are natural. We use ‘hierarchy’ notation to manage the constants we use, where $\alpha\ll\beta$ means that there is some non-negative decreasing function $f$ such that what follows will hold for all $\alpha\leq f(\beta)$. For hierarchies with more than 2 variables these functions are chosen from right to left. For any $a,b,c\in\mathbb{R}$, we say $a=b\pm c$ if $b-c\leq a\leq b+c$. We say an event occurs with high probability if it occurs with probability 1o(1)1-o(1).
@@ 2.2.1 Coupling processes
While doing this, we will simultaneously build a random graph $G_{*}$, starting similarly with the graph with vertex set $[n]$ and no edges. When we consider the random edge $e$, we also consider whether to add the random edge $e$ to $G_{*}$ or not; when $e\notin E(G_{*})$ we do this with the fixed probability $1-\eta$, where $0<\eta\ll\varepsilon$. Then, if $e$ is not in $F$ when it is considered and the probability at (1) is at least $1-\eta$, we can couple the random addition of $e$ so that $e$ is always added to $F$ if it is added to $G_{*}$. As discussed below, with high probability, for most of this process, the probability at (1) is very close to 1. Thus, if we build such a random graph $G_{*}$ and stop when it has $\lceil(1-2\eta)dn/2\rceil$ edges, then with high probability we will have $G_{*}\subset F$.
@@
This, however, is not good enough and our ‘error term’ $\eta$ in (4) will still need to change during the process. For some $C_{0},\mu$ with $1/n\ll 1/C_{0}\ll\mu\ll\varepsilon$, we will set $\eta_{m}=C_{0}\max\Big\{\frac{\mu}{\log n},\left(\frac{n\log n}{m}\right)^{1/8}\Big\}$, for each $m$, and our desired value of $\eta$ in (4) will be (roughly) $1-\eta_{e(F)}$, so that we can show that (3) is always at least the right hand side of (5) with (roughly) $\eta_{e(F)}$ replacing $\eta$. From (5) (and as proved in Section 3.5.5), as we will use this for each value of $e(F)$ from $\binom{n}{2}$ down to $\left\lfloor\left(1+\frac{\varepsilon}{2}\right)\frac{dn}{2}\right\rfloor$, one key point is that the sum of $\frac{\eta_{m}dn/2}{m-dn/2}$ over $m=e(F)$ in this range is $O(C_{0}\mu/\varepsilon)$, so that this will be much smaller than $\varepsilon$.
@@
Similar to Lemma 2.2, we will prove Theorem 2.4 using switching methods, but this will be significantly more difficult. Why this is more challenging is explained below, but the new ideas we use to do this are the heart of our ability to analyse this process as long as $d=\omega(\log n)$, compared to the regime $d\gg n/\sqrt{\log n}$ in which this was done by Gao, Isaev and McKay [15]. (In [13] a modified two-step process is used to neatly avoid some of these difficulties.)
@@
To use the switching method, we define a notion of ‘switching’ and consider when we can transform $K\in\mathcal{K}^{e}$ into $K^{\prime}\in\mathcal{K}^{f}$ using a switching. Comparing the number of ways we can do such a transformation on $K$ with the number of ways we could transform $K^{\prime}$, will allow us to compare $|\mathcal{K}^{e}|$ to $|\mathcal{K}^{f}|$. This is best seen via an example.
@@
and, as $\eta\ll 1$, we thus have $|\mathcal{K}^{e}|\leq(1+\eta)|\mathcal{K}^{f}|$.
Letting $K\in\mathcal{K}^{e}$, for each $K^{\prime}\in\mathcal{K}^{f}$ with $KK^{\prime}\in E(L)$, $E(K)\triangle E(K^{\prime})$ is the edge set of a length 10 cycle $S$ which alternates between $K\setminus F$ and $K_{n}\setminus K$ with $e$ and $f$ on opposite sides. Thus, we can bound $d_{L}(K)$ by counting such cycles. We will have that the vertex degrees of $K\setminus F$ are all $(1\pm\eta/100)\Delta$ for some $\Delta$ (in Lemma 2.2 this will be $2m/n$). Arbitrarily labelling $V(e)=\{v_{1},v_{6}\}$, we will use the notation in Figure 1 to choose such a cycle $S$ vertex by vertex. After labelling the vertices of $f$ as $v_{5}$ and $v_{10}$ (with 2 choices in total), and then choosing $v_{2}$ and $v_{7}$ (with at most $n$ choices for each), we then have at most $(1+\eta/100)\Delta$ choices for choosing each of $v_{3},v_{4},v_{8},v_{9}$, as they are the neighbour in $K\setminus F$ of some previously chosen vertex. Therefore, we have $d_{L}(K)\leq(1+\eta/100)^{4}\Delta^{4}n^{2}\leq(1+\eta/3)\Delta^{4}n^{2}$.
@@
For Lemma 2.2, as $F\sim F^{-}(n,d,m)$, with $m=\omega(n\log n)$ and $d=\omega(\log n)$, a simple concentration inequality (see Section 2.3) and a union bound shows that $K\setminus F$ will be approximately regular, allowing this sketch to be carried out when $d=o(n)$. For larger $d$ it is harder to count these switching cycles as accurately, as no longer is $K_{n}\setminus K$ so dense. In previous work [13], inequalities like (2) in such a regime were shown to be likely to hold using complex analytic methods. Here, we will use instead two simple ideas to complete the proof when $d=\Omega(n)$. These ideas are very useful for the upper part of the sandwich, so we will introduce them now in the simpler setting of the lower part.
@@
Then, in order to show it is likely that most of the graphs $K\in\mathcal{K}_{e}\cup\mathcal{K}_{f}$ satisfy some pseudorandom conditions, we exploit the distribution of $F\sim F^{-}(n,d,m)$. Intuitively, if we form $F$ by taking $K\sim G_{d}(n)$ and deleting $m$ random edges, then if a property is very likely to hold for $K$ it should be likely to hold for most of the $d$-regular graphs containing $F$. (This is proved with the exact parameters we use in Sections 9.1 and 10.5.) This means that instead of studying the properties of the $d$-regular graphs containing $F$, we can more simply study a random $d$-regular graph.
Switching methods for the upper part of the sandwich, i.e., for Theorem 2.4. After using the two ideas above, carrying out a similar switching argument for the upper part (i.e., for Theorem 2.4) will essentially require us to show that if $K\sim G_{d}(n)$ and $F$ is formed from $K$ by adding $m$ edges uniformly at random from $n]^{(2)}\setminus E(K)$ (and thus $F\sim F(n,d,m)$) then with high probability we have the following for some suitable $\ell$.
@@
Compared to similar tasks in the lower part, this is much more difficult in the region $d=o(n)$ as no longer is one of the $K$ or $F\setminus K$ graphs very close to being complete. Our previous discussion would correspond to paths of length $2\ell=4$, but now we will have to take $\ell=\Omega(\log n/\log\log n)$ when $d=\log^{O(1)}n$, otherwise between some distinct $x,y\in[n]$ no such alternating paths will exist. This pinpoints why the process for the upper part is so much harder to analyse via switchings than the process for the lower part. In their work, Gao, Isaev, and McKay [13] used two processes, one run after the other, so that (very roughly), when such $(F\setminus K,K)$-alternating paths were found, at least $K$ was sparse when compared to $F\setminus K$ (see [13] for more details).
When $F\setminus K$ is sufficiently dense, with $m\geq n^{1+\sigma}$ for any fixed $\sigma>0$, we use the results of Kim and Vu [19] on the concentration of multivariate polynomials to get sufficiently good bounds on the counts for A (see Section 8.2). Thus, we can fix some small $\sigma>0$, and assume that $d\leq n^{\sigma}$ and $m\leq n^{1+\sigma}$. Take distinct $x,y\in[n]$ and let $\delta=2m/n$, and note that we may expect $F\setminus K$ to have vertex degrees $(1+o(1))\delta$. Due to the randomness of $F\setminus K$ and $K$, if we randomly walk out from $x$ using edges alternately from $F\setminus K$ and then $K$ we should expect this random walk to rapidly mix so that, as long as (say) $\ell\geq 2\log n/\log(\delta d)$, after $2\ell$ edges we are roughly no more likely to arrive at $y$ than any other vertex. Heuristically, considering the vertex degrees in $F\setminus K$ and in the $d$-regular graph $K$, we should expect there are likely to be $(1+o(1))\delta^{\ell}d^{\ell}/n$ $(F\setminus K,K)$-alternating $x,y$-paths of length $2\ell$.
Bounds of this form were proved in certain regimes (roughly when $\delta\gg d\log n$) by Gao, Isaev and McKay [13] by working via the spectral properties of binomial random graphs (and using only that $K$ is $d$-regular), while similar counts on paths within a single graph have been shown by showing certain Markov chains mix rapidly (see, for example, as in [18], building on work in [8]). For Theorem 2.4, we need to work with the two different random graph models, in particular using properties of the more difficult to study random regular graph, and we need to deal with various subtleties due to the accuracy we need and the weaker bound on $d$ that we use. As far as is convenient, then, we rely on pseudorandom conditions of the random graphs involved, in part so that the results we prove may ultimately perhaps prove useful for the study of Conjecture 1.3.
@@
Roughly, then, we count walks alternating between $F\setminus K$ and $K$ using these sums of degrees, and show that most of these walks are actually paths. To do so we rely on pseudorandom properties that later appear as LABEL:eq:fewdoubled:K and LABEL:eq:fewdoubled:F-K in Lemma 6.3. Then, considering alternating paths from $x$ and $y$ respectively to sets of different endvertices we use lower bounds on the number of edges in $F\setminus K$ between these different sets of endvertices (using LABEL:prop:F-Kconnection later) to give a lower bound between the number of alternating paths between $x$ and $y$. A subtlety that is hard to anticipate from our discussion so far is that perhaps the most difficult case for Theorem 2.4 is when $d=C\log n$ and $\delta=\log^{O(1)}n$, but $\delta=\omega(\log^{8}n)$. This is because the error term, essentially $\eta$, that we wish to use in the counting of our alternating paths is to be smaller than $C_{0}\mu/\log n$ (see (6)), but we still have $\ell=\Omega(\log n/\log\log n)$, so that when we build our alternating paths of length $\ell$ from $x$, if the next edge from the current endpoint, say $v$, is to be in $K$ then $v$ may have a fraction of neighbours in the path so far which exceeds $\eta d$, which here is $O(1)$. This creates some additional small but important complication that, for example, is reflected in the error term we use in LABEL:prop:lowerboundKFKalternating:newnewnew in Theorem 6.1 later.
To show that the number of $(F\setminus K,K)$-alternating $x,y$-paths of length $2\ell$ is likely bounded above, we exploit our likely lower bound (applied with a slightly different path length) as well as an inductive upper bound on the number of alternating paths from $y$. Revealing the edges in $F\setminus K$ not containing $x$, we deduce that most vertices in $n]\setminus\{x,y\}$ – those not in $Z$, say – have at most $(1+O(\eta))\delta^{\ell-1}d^{\ell}/n$ $(K,F\setminus K)$-alternating paths of length $2\ell-1$ to $y$. Then, revealing the edges in $F\setminus K$ next to $x$, we get that there are at most $(1+O(\eta))\delta^{\ell}d^{\ell}/n$ $(F\setminus K,K)$-alternating $x,y$-paths of length $2\ell$ which do not have their second vertex in $Z\cap N_{F\setminus K}(x)$. As $Z$ was a relatively small set (it will have $O(\eta n)$ vertices), using this and similar properties iteratively will allow us to give an upper bound for all the $x,y$-paths we are interested in apart from those in an increasingly constrained set of paths. After sufficiently many iterations, a simple bound on the remaining possibilities will be sufficient to give us our desired upper bound. This argument is carried out in Section 7.
@@ 2.2.3 Paper outline
In Section 3, we give the details of the edge deletion process we use to create the upper part of the sandwich, reducing the proof of the upper part of the sandwich to Theorem 2.4. In Section 4, using our main switching argument we reduce proving Theorem 2.4 to proving certain pseudorandom conditions concerning switching paths are likely to hold for random graphs $F$ and $K$. In Sections 57 we prove these pseudorandom conditions are likely to hold when $m\leq n^{1+\sigma}$ (for any fixed $\sigma>0$) In Section 5, we give likely upper bounds on the number of $(F\setminus K,K)$-alternating paths of a given length which start from a fixed vertex. In Section 6, we give likely lower bounds on the number of $(F\setminus K,K)$-alternating paths between any two fixed vertices (moreover, on such paths which avoid an arbitrary small set of vertices). In Section 7, we give likely upper bounds on the number of $(F\setminus K,K)$-alternating paths between any two fixed vertices. In Section 8, we give similar bounds to Sections 57 when $m\geq n^{1+\sigma}$ (for any fixed small $\sigma$). In Section 9, we put this together to prove Theorem 2.4. In combination, this completes the proof for the upper part of the sandwich. In Section 10, we give the lower part of the sandwich and its analysis.